OpenStudy (anonymous):

the point (-3/5, y) in the third quadrant corresponds to angle θ on the unit circle. what is the value of sec θ

OpenStudy (gorv):



OpenStudy (gorv):

unit circle means x^2+y^2=1^2=1 x^2+y^2=1

OpenStudy (gorv):

we have (-3/5,y) this point on circle can u tell me the value of y after we put x=-3/5 in above equation??

OpenStudy (gorv):

(-3/5)^2+y^2=1 from here calculate y from here 2 values of y will come but we will consider value with negative side bcoozz in 3rd quad y is negative

OpenStudy (gorv):

you had x and y now calculate sec theta

OpenStudy (gorv):



OpenStudy (gorv):

use property of right angle triangle now ..and try ..u will surely get a ans